Lakisha owes her mechanic $1200 for a car repair. If he charges her 0.8% simple interest per month, and she makes no payments the first year, how much interest will she owe after the first year?
Answer by mananth(16946)

0.8% per month = 12*0.8 % per year
=9.6% per year.
..
I = p*r*t
I=1200*0.096*1
=$ 115.2 at the end of the year.